OpenStudy (anonymous):

does sin(sin(sin(sin(sin(...sin(x))))))))))))))....) converge to any number?

OpenStudy (anonymous):

it should go to zero

OpenStudy (anonymous):

0 i expect

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

because x can be anything and it will go from -1 to 1 and then from -1 to 1 u will get 0 eventually

OpenStudy (anonymous):

right

OpenStudy (anonymous):

but it may be \(\pm0\) :DD

OpenStudy (anonymous):

http://ideone.com/Yrttc

OpenStudy (anonymous):

http://ideone.com/TsQ7N

OpenStudy (anonymous):

but I'm sure there's a way to show this

OpenStudy (anonymous):

analytically

OpenStudy (anonymous):

Write out its sum: \[\sum_{k=0}^{\infty}\frac{(-1)^{2k+1}(\frac{\sum_{k=0}^{\infty} (-1)^{2k+1}(\frac{\sum_{k=0}^{\infty}...}{(2k+1)!})}{(2k+1)!})^{2k+1}}{(2k+1)!}\] And find the nth partial sum and check for convergence :D

OpenStudy (anonymous):

McLaurin series, whatever.

OpenStudy (anonymous):

it was question with sin, where did u get all this garbage? :D

OpenStudy (turingtest):

that is the mclauron series for sinx

OpenStudy (anonymous):

Math Education:\[\sin(x)=\sum_{k=0}^{\infty}\frac{(-1)^{2k+1}x^{2k+1}}{(2k+1)!}\]

OpenStudy (turingtest):

iterated many times... not sure that's solveable.

OpenStudy (anonymous):

Turing, just btw, I love your picture. And I've been reading your posts and you seem intelligent so whats up :D When I quit getting on here often you weren't around I think haha

OpenStudy (anonymous):

i know only geometric and arithmetic series :/

OpenStudy (anonymous):

It depends on how many times you compose sin(x) with itself right? n times or infinite? That changes the problem significantly.

OpenStudy (turingtest):

Yeah thanks malevolence. You definitely know your stuff better than me I think ;)

OpenStudy (anonymous):

i would try this. assuming it converges to say "y" then solve \[y=\sin(y)\]

OpenStudy (anonymous):

that gimmick presupposes that it converges to something.

OpenStudy (anonymous):

So you assume it converges. But the only thing you can do then is to take the arcsin over and over and over. So you get: \[\cos^{-1}(\cos^{-1}(...\cos^{-1}(L)))...)\] Assuming your limit is L. And Turing, probably not haha. I just try to be creative D:

OpenStudy (anonymous):

Those should be sin's***

OpenStudy (anonymous):

Oh check one more thing: look at that last post. http://openstudy.com/users/malevolence19#/updates/4ef73f34e4b01ad20b50e6fb

OpenStudy (turingtest):

Haha That's a heck of a rant there malevolence XD

OpenStudy (turingtest):

sounds like William S Burroughs...

OpenStudy (turingtest):

oh Steve Wright I see.

OpenStudy (anonymous):

I was just pulling stuff off the top of the head haha. And that integral didn't yield 0. x.x So apparently I set it up wrong or something. I got 4 pi. The thing is, I did the integral correctly because I got the same result by applying stokes theorem (vector calc). Idk /shrug/ what can I do?

OpenStudy (anonymous):

\[\Huge \oint\]

OpenStudy (turingtest):

Man, I'll have to pick that apart later, I just got done giving myself a headache trying to figure out what I did wrong on ruhmshalle's post :/

OpenStudy (anonymous):

Need help?

OpenStudy (anonymous):

It is clear based on satellite's argument that if it converges, then it must converge to zero. Assume x is in the interval [0, pi/2]. We can do this without loss of generality because if x isn't, then +/- sin(x) definitely is, so it quickly becomes irrelevant. Noting that sin(x) < x for all values of x and that y < x implies that sin(y) < sin(x), it is apparent that the sequence defined by \[a_n = \sin(a_{n-1}), a_0 = x \] is monotonically decreasing and has a liminf of zero. So, it converges to zero.

OpenStudy (anonymous):

Very, very slowly though...

OpenStudy (anonymous):

I should actually rephrase. It is not immediately clear that the liminf is zero (though I don't know what else it would be) but it is clear that the function is bounded by zero, and all monotonic bounded sequence converge. Satellite's argument takes over from there.

OpenStudy (anonymous):

sequences*

OpenStudy (anonymous):

I think you're looking at this with the wrong angle, there's a very simple approach using sequences... You have a sequence in the form \[U_{n}=f(U_{n-1})\] With f=sin This sequence will converge iff there's a point such that x=f(x), and clearly the only point that satisfies this is 0, so it converges on 0, QED.

OpenStudy (anonymous):

That is not always the case... if f is x^2, for instance, it actually diverges the other way. Most of us are not terribly familiar with fixed point iterations anyway.

OpenStudy (anonymous):

It works because sin is a strictly increasing function in [-pi;pi], and after the first iteration your term will obviously be in this interval.

OpenStudy (anonymous):

right. I want to find out what happens when n is infinity with that sequence someone wrote up there.

OpenStudy (anonymous):

There are more conditions that are necessary... x^2 is strictly increasing from 0 to infinity, and after the first iteration our term will be in that interval too.

OpenStudy (anonymous):

I'm not trying to be belligerent, I'm just trying to say that there are definitely conditions that need to be satisfied for the convergence of a floating point iteration to work out properly, and what I said above about the sequence being monotonically decreasing and bounded is sufficient. Maybe not using the language that you're used to, if you're familiar with F.P.I.'s, but it works.

OpenStudy (anonymous):

What are FPIs? Sorry, I'm not familiar with math in english

OpenStudy (anonymous):

I just abbreviated Fixed Point Iteration. And oops, I meant fixed point, not floating point o.O

OpenStudy (anonymous):

True, anyhow, it can be proved through the method I stated, I didn't type it out rigorously enough though... I think you can also do it with constantly decreasing intervals, since sin goes from R into [-1;1] and sin([-1;1]) will be an even smaller interval, then by iterations your intervals will become smaller and 0 will always be in them, meaning that for all nth iteration there will exist an interval that's a neighborhood of 0, meaning your sequence will tend to 0